OpenStudy (anonymous):

Glider A, traveling at 10 m/s on an air track, collides elastically with glider B traveling at 8.0 m/s in the same direction. The gliders are of equal mass. The final speed of glider B is (a) 8.4 m/s (b) 10 m/s (c) 8.0 m/s (d) 4.0 m/s (e) 12 m/s The answer is B, but I'm not sure why.

OpenStudy (michele_laino):

Please you have to apply the conservation momentum law and the conservation kinetic energy law

OpenStudy (shamim):

Let m1=m2=m

OpenStudy (shamim):

According to conservation of momentum m1*u1+m2*u2=m1*v1+m2*v2 m*10+m*8=m*v1+m*v2 18m=m1*v1+m2*v2.........(1)

OpenStudy (shamim):

Now u hv to use the law of conservation of kinetic energy (1/2)*m1*u1^2+(1/2)*m2*u2^2=(1/2)*m1*v1^2+(1/2)*m2*v2^2

OpenStudy (shamim):

50m+32m=.5m*v1^2+.5*m*v2^2............(2)

OpenStudy (shamim):

Frm equation (1) and (2) Find v1=? v2=?

OpenStudy (shamim):

Response plz!!!!